How do find the x and y intercept of 5x-6y= -15, and could you show me step by step?

Answer by Nate(3500)

To find the x-intercept, you plug 0 in for the y value.
5x-6y= -15 (original equation)
5x=-15 (0 times -6 gives you 0)
x=-3 (divide by 5)
the x-intercept is -3
To find the y-intercept, you plug 0 in for the x value.
5x-6y= -15 (original equation)
-6y=-15 (0 times 5 gives you 0)
y=15/6 (divide by -6)
the y-intercept is 15/6